# MERaLiON: Enhancing Cross-Lingual Understanding of Chinese, Indonesian, Malay and Singlish in Language Models

### v2

@@ 1 Introduction
Our first released model is named MERaLiON-LLaMA-3-8B-Instruct. MERaLiON-LLaMA-3-8B-Instruct has been extensively pre-trained on English, Chinese, and Indonesian, building upon LLaMA-3.1-8B-Base Dubey et al. (2024), with a primary emphasis on enhancing its understanding and generation capabilities in Chinese and Indonesian. Leveraging innovative corpus mixing strategies tailored to multilingual regional datasets, we carefully diversified the training materials using domain classification, hyperparameter optimization, and strategic replay techniques. These methods are specifically designed to prevent catastrophic forgetting, enabling the model to retain previously acquired knowledge while significantly improving its ability to produce high-quality, contextually accurate responses in Southeast Asian languages.
@@ 2 Pre-Training
MERaLiON-LLaMA-3-8B-Instruct training was conducted using the MaxText AI-Hypercomputer (2024) platform, utilizing both NVIDIA H100 GPUs and TPU v4-128 chips. Specifically, we utilized 64 H100 GPUs, achieving approximately 400 TFLOPS per GPU, and TPU v4-128 configurations, attaining around 168 TFLOPS per TPU chip.
@@ 3 Instruction Tuning and Model Merging
Model Cross-MMLU Cross-LogiQA English Chinese Indonesian English Chinese Indonesian MERaLiON-LLaMA-3-8B-Instruct 0.85 0.69 0.71 0.591 0.528 0.494 Meta-Llama-3.1-8B-Instruct 0.82 0.63 0.66 0.585 0.585 0.455 Instruction-Tuned Llama-3.1-8B 0.73 0.60 0.58 0.563 0.523 0.466
@@ 4 Benchmarks and Evaluation
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 MERaLiON-LLaMA-3-8B-Instruct 0.847 0.693 0.713 0.613 0.717 Meta-Llama-3.1-8B-Instruct 0.82 0.633 0.66 0.647 0.690 Llama3-8B-SEA-LION-v2.1-Instruct Singapore (2024) 0.753 0.667 0.693 0.64 0.688 Meta-Llama-3-8B-Instruct 0.767 0.653 0.573 0.573 0.642 Others GPT4o-0513 0.927 0.887 0.88 0.907 0.900 Gemma-2-9B-IT Team et al. (2024) 0.84 0.793 0.78 0.747 0.790 Gemma2-9B-SEA-LION-v3-Instruct Singapore (2024) 0.847 0.787 0.793 0.733 0.790 Qwen2.5-7B-Instruct Yang et al. (2024) 0.847 0.84 0.753 0.713 0.788 SeaLLMs-v3-7B-Chat Zhang et al. (2024) 0.833 0.727 0.74 0.687 0.747
@@
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 Meta-Llama-3.1-8B-Instruct 0.585 0.585 0.455 0.523 0.537 MERaLiON-LLaMA-3-8B-Instruct 0.591 0.528 0.494 0.489 0.526 Meta-Llama-3-8B-Instruct 0.602 0.523 0.438 0.483 0.512 Llama3-8B-SEA-LION-v2.1-Instruct 0.528 0.517 0.403 0.443 0.473 Others Qwen2.5-7B-Instruct 0.693 0.71 0.631 0.534 0.642 Gemma-2-9B-IT 0.659 0.636 0.585 0.602 0.621 Gemma2-9B-SEA-LION-v3-Instruct 0.636 0.642 0.557 0.551 0.597 SeaLLMs-v3-7B-Chat 0.568 0.585 0.494 0.517 0.541
@@ 4.1 Results
Model Series Model Accuracy LLaMA Series MERaLiON-LLaMA-3-8B-Instruct 0.576 Llama3-8B-SEA-LION-v2.1-Instruct 0.560 Meta-Llama-3.1-8B-Instruct 0.548 Meta-Llama-3-8B-Instruct 0.521 Others GPT4o-0513 0.760 Gemma2-9B-SEA-LION-v3-Instruct 0.626 Gemma-2-9B-IT 0.621 Qwen2.5-7B-Instruct 0.582 SeaLLMs-v3-7B-Chat 0.541
@@
Model Series Model Accuracy LLaMA Series MERaLiON-LLaMA-3-8B-Instruct 0.514 Llama3-8B-SEA-LION-v2.1-Instruct 0.505 Llama3-8B-SEA-LION-v2-Instruct 0.495 Meta-Llama-3-8B-Instruct 0.467 Meta-Llama-3.1-8B-Instruct 0.457 Others Qwen2-7B-Instruct 0.829 GPT4o-0513 0.810 Qwen2.5-7B-Instruct 0.800 Gemma2-9B-SEA-LION-v3-Instruct 0.590 Gemma-2-9B-IT 0.581
@@ 5 Conclusion and Future Work
The MERaLiON-LLaMA-3-8B-Instruct model improves multilingual NLP for Indonesian and Chinese, by addressing cross-lingual challenges with effective pretraining and model merging.

### v3

@@ 1 Introduction
Our first released model is named LLaMA-3-MERaLiON-8B-Instruct. LLaMA-3-MERaLiON-8B-Instruct has been extensively pre-trained on English, Chinese, and Indonesian, building upon LLaMA-3.1-8B-Base Dubey et al. (2024), with a primary emphasis on enhancing its understanding and generation capabilities in Chinese and Indonesian. Leveraging innovative corpus mixing strategies tailored to multilingual regional datasets, we carefully diversified the training materials using domain classification, hyperparameter optimization, and strategic replay techniques. These methods are specifically designed to prevent catastrophic forgetting, enabling the model to retain previously acquired knowledge while significantly improving its ability to produce high-quality, contextually accurate responses in Southeast Asian languages.
@@ 2 Pre-Training
LLaMA-3-MERaLiON-8B-Instruct training was conducted using the MaxText AI-Hypercomputer (2024) platform, utilizing both NVIDIA H100 GPUs and TPU v4-128 chips. Specifically, we utilized 64 H100 GPUs, achieving approximately 400 TFLOPS per GPU, and TPU v4-128 configurations, attaining around 168 TFLOPS per TPU chip.
@@ 3 Instruction Tuning and Model Merging
Model Cross-MMLU Cross-LogiQA English Chinese Indonesian English Chinese Indonesian LLaMA-3-MERaLiON-8B-Instruct 0.85 0.69 0.71 0.591 0.528 0.494 Meta-Llama-3.1-8B-Instruct 0.82 0.63 0.66 0.585 0.585 0.455 Instruction-Tuned Llama-3.1-8B 0.73 0.60 0.58 0.563 0.523 0.466
@@ 4 Benchmarks and Evaluation
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 LLaMA-3-MERaLiON-8B-Instruct 0.847 0.693 0.713 0.613 0.717 Meta-Llama-3.1-8B-Instruct 0.82 0.633 0.66 0.647 0.690 Llama3-8B-SEA-LION-v2.1-Instruct Singapore (2024) 0.753 0.667 0.693 0.64 0.688 Meta-Llama-3-8B-Instruct 0.767 0.653 0.573 0.573 0.642 Others GPT4o-0513 0.927 0.887 0.88 0.907 0.900 Gemma-2-9B-IT Team et al. (2024) 0.84 0.793 0.78 0.747 0.790 Gemma2-9B-SEA-LION-v3-Instruct Singapore (2024) 0.847 0.787 0.793 0.733 0.790 Qwen2.5-7B-Instruct Yang et al. (2024) 0.847 0.84 0.753 0.713 0.788 SeaLLMs-v3-7B-Chat Zhang et al. (2024) 0.833 0.727 0.74 0.687 0.747
@@
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 Meta-Llama-3.1-8B-Instruct 0.585 0.585 0.455 0.523 0.537 LLaMA-3-MERaLiON-8B-Instruct 0.591 0.528 0.494 0.489 0.526 Meta-Llama-3-8B-Instruct 0.602 0.523 0.438 0.483 0.512 Llama3-8B-SEA-LION-v2.1-Instruct 0.528 0.517 0.403 0.443 0.473 Others Qwen2.5-7B-Instruct 0.693 0.71 0.631 0.534 0.642 Gemma-2-9B-IT 0.659 0.636 0.585 0.602 0.621 Gemma2-9B-SEA-LION-v3-Instruct 0.636 0.642 0.557 0.551 0.597 SeaLLMs-v3-7B-Chat 0.568 0.585 0.494 0.517 0.541
@@ 4.1 Results
Model Series Model Accuracy LLaMA Series LLaMA-3-MERaLiON-8B-Instruct 0.576 Llama3-8B-SEA-LION-v2.1-Instruct 0.560 Meta-Llama-3.1-8B-Instruct 0.548 Meta-Llama-3-8B-Instruct 0.521 Others GPT4o-0513 0.760 Gemma2-9B-SEA-LION-v3-Instruct 0.626 Gemma-2-9B-IT 0.621 Qwen2.5-7B-Instruct 0.582 SeaLLMs-v3-7B-Chat 0.541
@@
Model Series Model Accuracy LLaMA Series LLaMA-3-MERaLiON-8B-Instruct 0.514 Llama3-8B-SEA-LION-v2.1-Instruct 0.505 Llama3-8B-SEA-LION-v2-Instruct 0.495 Meta-Llama-3-8B-Instruct 0.467 Meta-Llama-3.1-8B-Instruct 0.457 Others Qwen2-7B-Instruct 0.829 GPT4o-0513 0.810 Qwen2.5-7B-Instruct 0.800 Gemma2-9B-SEA-LION-v3-Instruct 0.590 Gemma-2-9B-IT 0.581
@@ 5 Conclusion and Future Work
The LLaMA-3-MERaLiON-8B-Instruct model improves multilingual NLP for Indonesian and Chinese, by addressing cross-lingual challenges with effective pretraining and model merging.